OpenStudy (anonymous):

Use the Integral Test to determine whether the series is convergent or divergent

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty }n*e^{-12n}\]

OpenStudy (michele_laino):

please you can use the n-th roots criterion too, since your series is a series of positive terms

OpenStudy (anonymous):

what is that?

OpenStudy (michele_laino):

please note that if the subsequent quantity: \[\mathop {\lim }\limits_{n \to \infty } \sqrt[n]{{{a_n}}} = L\quad and\;L < 1\] then the positive terms series \[\sum\limits_n {{a_n}} \] is a convergent series

OpenStudy (anonymous):

is "e" a square root like that?

OpenStudy (michele_laino):

Now in your case we have: \[\begin{gathered} {a_n} = \frac{n}{{{e^{12n}}}} \hfill \\ \mathop {\lim }\limits_{n \to \infty } \sqrt[n]{{{a_n}}} = \mathop {\lim }\limits_{n \to \infty } \sqrt[n]{{\frac{n}{{{e^{12n}}}}}} = \mathop {\lim }\limits_{n \to \infty } \frac{{\sqrt[n]{n}}}{{{e^{12}}}} = \frac{1}{{{e^{12}}}} < 1 \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

oh wow.. thank you!

OpenStudy (michele_laino):

thank you!